# Dynamic and Thermodynamic Stability of Superconducting-superfluid Stars

### v1

@@ Abstract
We give a comprehensive analysis of the dynamic and thermodynamic stability of neutron stars composed of superconducting-superfluid mixtures within the Iyer-Wald formalism. We derive the first law of thermodynamics and the necessary and sufficient condition under which dynamic equilibrium implies thermodynamic equilibrium. By constructing the phase space and canonical energy, we show that the dynamic stability for perturbations, restricted in symplectic complement of trivial perturbations with the ADM 3-momentum unchanged, is equivalent to the non-negativity of the canonical energy. Furthermore, dynamic stability against restricted axisymmetric perturbations guarantees the dynamic stability against all axisymmetric perturbations. We also prove that the positivity of canonical energy on all axisymmetric perturbations within the Lagrangian displacement framework with fixed angular momentum is necessary for thermodynamic stability. In particular, the equivalence of dynamic and thermodynamic stability for spherically symmetric perturbations of static, spherically symmetric isentropic configurations is established.
@@ I Introduction
The dynamic stability of self-gravitating compact objects is a cornerstone of theoretical astrophysics, determining whether they end up as a stable stars, collapsing to black holes, or exploding in a supernova. For the relativistic neutral or charged stars described by the single perfect fluid model, the criterion for dynamic stability has been established, and it has been found to be closely related to the thermodynamic stability [Green:2013ica, Shi:2022jya]. Specifically, the criterion for dynamic stability of perfect fluid star in dynamic equilibrium is given by non-negativity of the canonical energy associated with the timelike Killing field, and the necessary condition for the thermodynamic stability of stars in thermodynamic equilibrium with respect to the axisymmetric perturbations is the positivity of the canonical energy. Furthermore, the dynamic and thermodynamic stability are equivalent if the background star is static, spherically symmetric isentropic and the perturbations are spherically symmetric.
Although the single perfect fluid model is a remarkably good approximate for many types of stars, such as the red giants and white dwarfs, it can only serve the roughest description of the “multi-constituent fluid stars”, in particular, the neutron stars. The neutron star, excluding an outer magnetosphere of negligible mass and an inner core consisting of matter in exotic state, is broadly understood that can be described in terms of three principle layers: the outer crust made of a lattice of nuclei with gas of relativistic degenerate electrons throughout the star, the inner crust whose solid ionic lattice is interpenetrated by a neutron superfluid, and the outer core consisting of a neutron superfluid and superconducting proton fluid [Langlois:2000gs]. In the relevant layers, the superfluid neutrons make up the most of the mass density, the superconducting protons make up a small but significant part of the mass density, and the degenerate non-superconducting electrons make up a negligibly small fraction of the mass density, but nevertheless have an important role when the electromagnetic effects are concerned. Therefore, a more accurate approximation for the neutron star should be the multi-constituent fluid model.
@@
The two-fluid model of non-relativistic superfluid is developed by Laudau, and then generalized to the relativistic case and multi-constituent fluid model by Khalatnikov and Carter [KHALATNIKOV198270, Carter:1987qr, Carter:1992gmy, CARTER1992243, Carter:1993aq]. For a basic representation of the neutron star’s superconducting superfluid region, the two-fluid model is often adequate, where one constituent is the superfluid neutron while the other represents everything else, i.e., the approximately rigid background consists of protons and electrons that are tended by the short range electromagnetic interactions. Particularly, the two-constituent fluid model including allowance for “transfusion”, meaning the slow transfer of baryonic matter (due to the process like beta decay and so on) between the neutron superfluid and the other “normal” fluid, has been constructed by Langlois et al. [Langlois:1997bz]. However, because of its neglect of the electromagnetically interacting constituents, which will play an essential role in phenomena involving magnetic effects, the superconducting-superfluid mixtures [Carter:1998rn] would be a more suitable framework.
@@
Recently, there have been some studies on the radial stability of neutron stars [Caballero:2024qtv, Canullan-Pascual:2024jsu, Caballero:2025omv, Kumar:2025oyx]. For instance, reference [Caballero:2024qtv] establishes a set of stability criteria for two perfect-fluid relativistic star by studying the radial mode perturbation equations, and provides also an alternative stability criterion (i.e., the positivity of canonical energy) in the same way as done for single perfect fluid stars. In [Kumar:2025oyx], the eigenvalue problem for a coupled system of equations with small-amplitude radial perturbations is solved and the critical line corresponding to stability boundaries is derived. Nevertheless, they consider only two non-interacting perfect fluids, which are the simplification of the two-fluid model mentioned above where there are non-gravitational interactions between the two fluids in general, and the background star is taken to be static and spherically symmetric.
@@
In this paper, we consider the relativistic stars described by the non-transfusive superconducting-superfluid mixtures. By working with Iyer-Wald formalism [Lee:1990nz, Wald:1993nt, Iyer:1994ys] rather than analyzing the perturbation equations, we will give the necessary and sufficient conditions for thermodynamic equilibrium, construct the phase space of our system, and then establish the criteria for both dynamic stability and thermodynamic stability as the positivity of canonical energy, which are equivalent if the background star is static, spherically symmetric isentropic and the perturbations are spherically symmetric, i.e., the radial perturbations. In our stability analysis, which includes the treatment of trivial displacements, we employ the most general form of Eulerian perturbations without any gauge fixing. This approach is necessitated by the presence of multiple Lagrangian displacements in a multi-constituent system and is more general than the single perfect fluid case, where a simplification to the Lagrangian perturbations, such as the gauge used in [Green:2013ica, Shi:2022jya], is possible.
@@
Similar to the arguments given in the introduction of [Green:2013ica], although our approach is completely incapable of yielding any information concerning the growth rate of any thermodynamic instability, it should lead to results equivalent to those that obtained by considering dissipation in principle. Furthermore, our results can be directly generalized to the case of superconducting-superfluid mixtures with one “normal” component and arbitrary number of “super” components, and the case including the allowance for transfusion.
@@
The rest of this paper is structured as follows. In Sec. II, we review relativistic model of the superconducting-superfluid mixtures. In Sec. III, we give a quick introduction to the Iyer-Wald formalism, and list some results that will be used in later. In Sec. IV, we derive the first law of thermodynamics, and after defining the thermodynamic equilibrium, the necessary and sufficient condition for the star in thermodynamic equilibrium is derived. Although the fluid model we used does not include the transfusion, but we will also give a simple argument about the transfusive case and show one of our definitions is suitable for such case at the end of this section. In Sec. V, we devote ourselves to constructing the phase space and calculating the symplectic complement of trivial perturbations. With such preparations, the criterion for the dynamic stability is established by introducing the canonical energy and taking advantage of the physically stationary perturbations in Sec. VI, and the necessary condition for the thermodynamic stability is established in Sec. VII. Finally, we give the conclusion and discussion in Sec. VIII. In Appendix A we derive the eigenvalue equation of propagation velocity $v$ of the sound wave, which will be useful in finding the degeneracy of pre-symplectic form, by examining the characteristic hypersurfaces of discontinuity.
@@ II Relativistic model of superconducting-superfluid mixtures
In this section, we will review the Carter’s relativistic model of superconducting-superfluid mixtures [Carter:1998rn]. The three independent constituents under consideration are the superfluid neutrons with conserved particle current $n_{\text{n}}^{a}$, the superconducting protons with conserved particle current $n_{\text{p}}^{a}$, and the degenerate non-superconducting background of electrons with conserved particle current $n_{\text{e}}^{a}$. The electrons and protons have crucially important roles as electromagnetic effects are concerned, and in terms of the electron charge coupling constant $e$ the corresponding total electric current vector will be given by
@@
Besides three principal constituents that have just been listed, there is a fourth constituent, namely the conserved entropy current $s^{a}=sn_{\text{e}}^{a}$, which is carried by the “normal” electron fluid and $s$ is the entropy per electron. For convenience, we shall use the capital Latin letters $\text{X}=\text{n},\text{p},\text{e}$ below as the “chemical” indices of three relevant constituents, and using this convention, the equation Eq. (1) can be rewritten in the concise form
@@ II.1 Master function
The central quantity of Carter’s theory of multi-constituent fluid is the so-called master function [Carter:1987qr], which is taken to be the total thermodynamic energy density $\Lambda_{\text{M}}$, depending only on the metric $g_{ab}$ and the “hydrodynamic” part of the system, i.e., $n_{\text{n}}^{a}$, $n_{\text{p}}^{a}$, $n_{\text{e}}^{a}$ and $s$. More exactly, $\Lambda_{\text{M}}$ is a function of all scalar combinations obtained by their mutual contractions
@@
is the inertia matrix and we will assume that it is positive definite as it in the two fluid model [CARTER1992243].
@@ II.2 Dynamical fields, variations, and Lagrangian displacements
To develop a Lagrangian description of the superconducting-superfluid mixtures, not only are we required to have the spacetime manifold $\mathcal{M}$, on which the metric $g_{ab}$ and the electromagnetic potential $A_{a}$ are defined, but also for each constituent we should introduce a fiducial manifold $\mathcal{M}_{\text{X}}^{\prime}$, called fluid spacetime, which is diffeomorphic to $\mathcal{M}$. Then with a fixed scalar field $s^{\prime}$ on $\mathcal{M}_{\text{e}}^{\prime}$, and fixed 3-form $\bm{N}^{\text{X}\prime}$ on $\mathcal{M}_{\text{X}}^{\prime}$, one can define the physical fluid fields on $\mathcal{M}$ by pushing forward with diffeomorphism $\chi_{\text{X}}$ as
@@ II.3 Lagrangian of superconducting-superfluid mixtures
and the energy-momentum tensor is given by [^1] It is easy to check that $n_{\text{X}}^{a}\mu^{\text{X}b}=n_{\text{X}}^{b}\mu^{\text{X}a}$ by using Eq. (5).
@@
with the generalization of pressure [Carter:1987qr]
@@ III Lagrangian framework for diffeomorphism covariant theories
where $\mathcal{D}$ represents the exterior derivative on forms on $\mathcal{F}$. For 1-form $\Theta_{A}$, one has
@@
$d\bm{\omega}\left(\phi;\delta_{1}\phi,\delta_{2}\phi\right)=\delta_{2}\bm{E}\cdot\delta_{1}\phi-\delta_{1}\bm{E}\cdot\delta_{2}\phi,$ (37)
@@
For a diffeomorphism covariant Lagrangian, the Noether current 3-form on spacetime associated with an arbitrary vector field $X^{a}$ is defined by
@@
A simple calculation [Iyer:1994ys] shows that the first variation of $\bm{\mathcal{J}}_{X}$ (with $X^{a}$ fixed, i.e., unvaried) satisfies
$\delta\bm{\mathcal{J}}_{X}=-\iota_{X}\left(\bm{E}\cdot\delta\phi\right)+\bm{\omega}\left(\phi;\delta\phi,\mathscr{L}_{X}\phi\right)+d\left[\iota_{X}\bm{\theta}\left(\phi;\delta\phi\right)\right],$ (39)
where it has not been assumed that $\phi$ satisfies the field equations nor that $\delta\phi$ satisfies the linearized field equations. Furthermore, it can be shown that $\bm{\mathcal{J}}_{X}$ can be written in the form [Iyer:1995kg]
$\bm{\mathcal{J}}_{X}=\bm{C}_{X}+d\bm{Q}_{X},$ (40)
where $\bm{Q}_{X}$ is the Noether charge and $\bm{C}_{X}\equiv X^{a}\bm{C}_{a}$ with $\bm{C}_{a}=0$ being the constraint equations of the theory [Seifert:2006kv]. Having written in this form, we obtain the fundamental identity
$\bm{\omega}\left(\phi;\delta\phi,\mathscr{L}_{X}\phi\right)=\iota_{X}\bm{E}\cdot\delta\phi+\delta\bm{C}_{X}+d\left[\delta\bm{Q}_{X}-\iota_{X}\bm{\theta}\left(\phi;\delta\phi\right)\right],$ (41)
It should be emphasized that this fundamental identity holds for arbitrary $X^{a}$, $\phi$, and $\delta\phi$.
One immediate consequence of Eq. (41) is the gauge invariance of the symplectic form. If $\phi$ satisfies the equations of motion, $\bm{E}=0$, $\delta\phi$ satisfies the linearized constraints, $\delta\bm{C}_{a}=0$, and $X^{a}$ is of compact support (or vanishes sufficiently rapidly at infinity and/or any boundaries), integration of Eq. (41) over a Cauchy surface $\Sigma$ yields
$W_{AB}\delta\phi^{A}\mathscr{L}_{X}\phi^{B}=0.$ (42)
Consequently, the value of $W_{AB}\delta\phi^{A}\delta\phi^{B}$ is unchanged if either $\delta_{1}\phi$ or $\delta_{2}\phi$ is altered by a gauge transformation $\delta\phi\rightarrow\delta\phi+\mathscr{L}_{X}\phi$ with $X^{a}$ of compact support.
Another very important application concerns the case where $X^{a}$ approaches a nontrivial asymptotic symmetry rather than being of compact support, in which case we can derive a formula for the Hamiltonian, $H_{X}$, conjugate to the notion of “translations” defined by $X^{a}$, and, thereby, a definition of ADM-type conserved quantities. Consider asymptotically flat spacetime with one asymptotically flat “end”. Integrating Eq. (41) over a Cauchy surface $\Sigma$, we have
$W_{AB}\delta\phi^{A}\mathcal{L}_{X}\phi^{B}$
$=$ $\int_{\Sigma}\left(\iota_{X}\bm{E}\cdot\delta\phi+\delta\bm{C}_{X}\right)+\int_{S_{\infty}}\left[\delta\bm{Q}_{X}-\iota_{X}\bm{\theta}\left(\phi;\delta\phi\right)\right],$ (43)
@@
$\lim_{S\rightarrow S_{\infty}}\int_{S}\iota_{X}\bm{\theta}\left(\phi;\delta\phi\right)=\lim_{S\rightarrow S_{\infty}}\delta\int_{S}\iota_{X}\bm{B},$ (44)
@@
$W_{AB}\delta\phi^{A}\mathscr{L}_{X}\phi^{B}=\delta H_{X},$ (45)
@@
$H_{X}=\int_{\Sigma}\bm{C}_{X}+\int_{S_{\infty}}\left(\bm{Q}_{X}-\iota_{X}\bm{B}\right).$ (46)
@@
$\delta H_{X}=\delta\phi^{A}D_{A}H_{X}=\left(\mathcal{D}H_{X}\right)_{A}\delta\phi^{A},$ (47)
@@
$W_{AB}\mathscr{L}_{X}\phi^{B}=\left(\mathcal{D}H_{X}\right)_{A}.$ (48)
@@
$\left(\mathscr{L}_{X}\phi\right)^{A}=W^{AB}\left(\mathcal{D}H_{X}\right)_{B},$ (49)
which is the usual form of Hamilton’s equations of motion on a symplectic manifold. Thus if both the asymptotic conditions on $\phi$ and the asymptotic behavior of $X^{a}$ are such that a 3-form $\bm{B}$ satisfying Eq. (44) exists, then Eq. (46) yields a Hamiltonian conjugate to the notion of “translations” defined by $X^{a}$. Note that when evaluated on solutions, $\bm{C}_{X}=0$, $H_{X}$ is purely a “surface term”
$H_{X}|_{\bm{E}=0}=\int_{S_{\infty}}\left(\bm{Q}_{X}-\iota_{X}\bm{B}\right).$ (50)
In the case where $X^{a}$ is asymptotic to a time translation $t^{a}$ at infinity, and $\bm{B}$ satisfying Eq. (44) can be found, then Eq. (50) defines the ADM mass
@@
In the case where $X^{a}$ is asymptotic to a rotation $\varphi^{a}$ tangent to $\Sigma$ at infinity and $S$ is chosen so that $X^{a}$ is tangent to $S$, the pull back of $\iota_{X}\bm{\theta}$ to $S$ vanishes, then Eq. (50) with $X^{a}=\varphi^{a}$ and $\bm{B}=0$ defines minus the ADM angular momentum

### v2

@@ Abstract
We present a comprehensive analysis of the dynamic and thermodynamic stability of neutron stars composed of superconducting-superfluid mixtures within the Iyer-Wald formalism. We derive the first law of thermodynamics and the necessary and sufficient condition under which dynamic equilibrium implies thermodynamic equilibrium. By constructing the phase space and canonical energy, we show that the dynamic stability for perturbations, restricted in symplectic complement of trivial perturbations with the ADM 3-momentum unchanged, is equivalent to the non-negativity of the canonical energy. Furthermore, dynamic stability against restricted axisymmetric perturbations guarantees the dynamic stability against all axisymmetric perturbations. We also prove that the positivity of canonical energy on all axisymmetric perturbations within the Lagrangian displacement framework with fixed angular momentum is necessary for thermodynamic stability. In particular, the equivalence of dynamic and thermodynamic stability is established for radial perturbations of static, spherically symmetric isentropic configurations.
@@ I Introduction
The dynamic stability of self-gravitating compact objects is a cornerstone of theoretical astrophysics, determining whether they end up as a stable stars, collapsing to black holes, or exploding in a supernova. For the relativistic neutral or charged stars described by the single perfect fluid model, the criterion for dynamic stability has been established, and it has been found to be closely related to the thermodynamic stability [19, 32]. Specifically, the criterion for dynamic stability of perfect fluid star in dynamic equilibrium is given by non-negativity of the canonical energy associated with the timelike Killing field, and the necessary condition for the thermodynamic stability of stars in thermodynamic equilibrium with respect to the axisymmetric perturbations is the positivity of the canonical energy. Furthermore, the dynamic and thermodynamic stability are equivalent if the background star is static, spherically symmetric isentropic and the perturbations are spherically symmetric.
While the single perfect fluid model provides an excellent approximation for many stellar objects such as red giants and white dwarfs, it is fundamentally inadequate for neutron stars due to their exceptionally complex internal microphysics. The interior of a neutron star is a multi-constituent system exhibiting distinct states across its layers. Broadly, its structure comprises a solid crystalline crust, a liquid outer core, and a potentially exotic inner core, with the equation of state and composition being active areas of research constrained by multi-messenger observations [27, 30, 1, 3, 29]. Crucially for macroscopic dynamics, a defining feature of the relevant layers (the inner crust and the outer core) is the prevalence of superfluidity and superconductivity. In these regions, neutrons are believed to form a superfluid, while protons are in superconducting state, all permeated by a degenerate electron gas [20, 16]. In the relevant layers, the superfluid neutrons make up the most of the mass density, the superconducting protons make up a small but significant part of the mass density, and the degenerate non-superconducting electrons make up a negligibly small fraction of the mass density, but nevertheless have an important role when the electromagnetic effects are concerned. This multi-component nature is not a minor correction but the central paradigm for explaining key observational phenomena. Therefore, a physically realistic hydrodynamic description of neutron star interiors must move beyond the single-fluid approximation. A multi-constituent fluid model, accounting for the distinct dynamics between the superfluid neutrons, the superconducting protons, and the electron plasma, is essential.
@@
The two-fluid model of non-relativistic superfluid is developed by Laudau, and then generalized to the relativistic case and multi-constituent fluid model by Khalatnikov and Carter [24, 15, 10, 9, 11]. For a basic representation of the neutron star’s superconducting superfluid region, the two-fluid model is often adequate, where one constituent is the superfluid neutron while the other represents everything else, i.e., the approximately rigid background consists of protons and electrons that are tended by the short range electromagnetic interactions. Particularly, the two-constituent fluid model including allowance for “transfusion”, meaning the slow transfer of baryonic matter (due to the process like beta decay and so on) between the neutron superfluid and the other “normal” fluid, has been constructed by Langlois et al. [26]. However, because of its neglect of the electromagnetically interacting constituents, which will play an essential role in phenomena involving magnetic effects, the superconducting-superfluid mixtures [14] would be a more suitable framework.
@@
The stability properties of non-relativistic rotating superfluid neutron stars have been studied long ago by using the canonical energy formulation [2]. And more recently, there have been some studies on the radial stability of neutron stars in relativistic case [6, 8, 7, 25]. For instance, reference [6] establishes a set of stability criteria for two perfect-fluid relativistic star, which is taken as a model of dark matter admixed neutron stars, by studying the radial mode perturbation equations, and provides also an alternative stability criterion (i.e., the positivity of canonical energy) in the same way as done for single perfect fluid stars. In [25], the eigenvalue problem for a coupled system of equations with small-amplitude radial perturbations is solved and the critical line corresponding to stability boundaries is derived. Nevertheless, most of above researchers concern the model of two non-interacting perfect fluids, which are the simplification of the two-fluid model mentioned above where there are non-gravitational interactions between the two fluids in general, and the background star is taken to be spherically symmetric.
@@
In this paper, we consider the relativistic neutron stars described by the non-transfusive superconducting-superfluid mixtures. By working with Iyer-Wald formalism [28, 35, 22] rather than analyzing the perturbation equations, we will give the necessary and sufficient conditions for thermodynamic equilibrium, construct the phase space of our system, and then establish the criteria for both dynamic stability and thermodynamic stability as the positivity of canonical energy, which are equivalent if the background star is static, spherically symmetric isentropic and the perturbations are spherically symmetric, i.e., the radial perturbations. In our stability analysis, which includes the treatment of trivial displacements, we employ the most general form of Eulerian perturbations without any gauge fixing. This approach is necessitated by the presence of multiple Lagrangian displacements in a multi-constituent system and is more general than the single perfect fluid case, where a simplification to the Lagrangian perturbations, such as the gauge used in [19, 32], is possible.
@@
Similar to the arguments given in the introduction of [19], when some perturbations acting on the star, if it is thermodynamically stable, the dissipative processes will make it back to the equilibrium, otherwise the dissipative processes will make it away from equilibrium. Our criterion for thermodynamic stability is established by examining the behavior of the entropy $S$, and it should be equivalent to the criterion by examining the dissipative dynamics as the dissipative processes tend to drive entropy to increase. So although our approach is incapable of yielding any information concerning the growth rate of any thermodynamic instability, it should lead to results equivalent to those that obtained by considering dissipation in principle. Furthermore, our results can be directly generalized to the case of superconducting-superfluid mixtures with one “normal” component and arbitrary number of “super” components, and even the case including the allowance for transfusion.
@@
The rest of this paper is structured as follows. In Sec. II, we review relativistic model of the superconducting-superfluid mixtures. In Sec. III, we give a quick introduction to the Iyer-Wald formalism, and list some results that will be used in later. In Sec. IV, we derive the first law of thermodynamics, and after defining the thermodynamic equilibrium, the necessary and sufficient condition for the star in thermodynamic equilibrium is derived. Although the fluid model we used does not include the transfusion, but we will also give a simple argument about the transfusive case and show one of our definitions is suitable for such case at the end of this section. In Sec. V, we devote ourselves to constructing the phase space and calculating the symplectic complement of trivial perturbations. As a by-product, we also derive the eigenvalue equation of propagation velocity $v$ of the sound wave, which will be useful in finding the degeneracy of pre-symplectic form. With such preparations, the criterion for the dynamic stability is established by introducing the canonical energy and taking advantage of the physically stationary perturbations in Sec. VI, and the necessary condition for the thermodynamic stability is established in Sec. VII. Finally, we give the conclusion and discussion in Sec. VIII.
@@ II Relativistic model of superconducting-superfluid mixtures
In this section, we will review the Carter’s relativistic model of superconducting-superfluid mixtures [14]. The three independent constituents under consideration are the superfluid neutrons with conserved particle current $n_{\text{n}}^{a}$, the superconducting protons with conserved particle current $n_{\text{p}}^{a}$, and the degenerate non-superconducting background of electrons with conserved particle current $n_{\text{e}}^{a}$. The electrons and protons have crucially important roles as electromagnetic effects are concerned, and in terms of the electron charge coupling constant $e$ the corresponding total electric current vector will be given by
@@
Besides three principal constituents that have just been listed, there is a fourth constituent, namely the conserved entropy current $s^{a}=sn_{\text{e}}^{a}$. Within the standard multi-constituent fluid model for superfluids, the entropy is associated exclusively with the normal (non-superfluid) component. Accordingly, as we consider that all neutrons and protons are in “super” states in our model, the entropy current is only carried by the “normal” electron fluid with $s$ denoting the entropy per electron. For convenience, we shall use the capital Latin letters $\text{X}=\text{n},\text{p},\text{e}$ below as the “chemical” indices of three relevant constituents, and using this convention, the equation Eq. (1) can be rewritten in the concise form
@@ II.1 Master function
The central quantity of Carter’s theory of multi-constituent fluid is the so-called master function [15], which is taken to be the total thermodynamic energy density $\Lambda_{\text{M}}$, depending only on the metric $g_{ab}$ and the “hydrodynamic” part of the system, i.e., $n_{\text{n}}^{a}$, $n_{\text{p}}^{a}$, $n_{\text{e}}^{a}$ and $s$. More exactly, $\Lambda_{\text{M}}$ is a function of all scalar combinations obtained by their mutual contractions
@@
is the inertia matrix and we will assume that it is positive definite like it in the two fluid model [9].
@@ II.2 Dynamical fields, variations, and Lagrangian displacements
To develop a Lagrangian description of the superconducting-superfluid mixtures, not only are we required to have the spacetime manifold $\mathcal{M}$, on which the metric $g_{ab}$ and the electromagnetic potential $A_{a}$ are defined, but also for each constituent we should introduce a fiducial manifold $\mathcal{M}_{\text{X}}^{\prime}$, called fluid spacetime, which is diffeomorphic to $\mathcal{M}$. Then with a fixed scalar field $s^{\prime}$ on $\mathcal{M}_{\text{e}}^{\prime}$, and fixed closed 3-form $\bm{N}^{\text{X}\prime}$ on $\mathcal{M}_{\text{X}}^{\prime}$, one can define the physical fluid fields on $\mathcal{M}$ by pushing forward with diffeomorphism $\chi_{\text{X}}$ as
@@ II.3 Lagrangian of superconducting-superfluid mixtures
and the energy-momentum tensor is given by [^1] It is easy to check that $\sum_{\text{X}}n_{\text{X}}^{a}\mu^{\text{X}b}=\sum_{\text{X}}n_{\text{X}}^{b}\mu^{\text{X}a}$ by using Eq. (5).
@@
with the generalization of pressure [15]
@@ III Lagrangian framework for diffeomorphism covariant theories
where $\mathcal{D}$ represents the exterior derivative on forms on $\mathcal{F}$. For 1-form $\Theta_{A}$, one has [^2] A standard result of differential geometry states that for 1-form $\bm{\mathcal{Y}}$ and vectors $\mathcal{X}_{1},\mathcal{X}_{2}$, $d\bm{\mathcal{Y}}\left(\mathcal{X}_{1},\mathcal{X}_{2}\right)=\mathcal{X}_{1}\left(\bm{\mathcal{Y}}\left(\mathcal{X}_{2}\right)\right)-\mathcal{X}_{2}\left(\bm{\mathcal{Y}}\left(\mathcal{X}_{1}\right)\right)-\bm{\mathcal{Y}}\left(\left[\mathcal{X}_{1},\mathcal{X}_{2}\right]\right).$ (32)
@@
$d\bm{\omega}\left(\phi;\delta_{1}\phi,\delta_{2}\phi\right)$
$=$ $\delta_{1}d\bm{\theta}\left(\phi;\delta_{2}\phi\right)-\delta_{2}d\bm{\theta}\left(\phi;\delta_{1}\phi\right)-d\bm{\theta}\left(\phi;\delta_{1}\delta_{2}\phi-\delta_{2}\delta_{1}\phi\right)$
[…]
$=$ $\delta_{2}\bm{E}\cdot\delta_{1}\phi-\delta_{1}\bm{E}\cdot\delta_{2}\phi,$ (38)
@@
For a diffeomorphism covariant Lagrangian, the Noether current 3-form on spacetime associated with an arbitrary vector field $\mathcal{X}^{a}$ is defined by
@@
A simple calculation [22] shows that the first variation of $\bm{\mathcal{J}}_{\mathcal{X}}$ (with $\mathcal{X}^{a}$ fixed, i.e., unvaried) satisfies
$\delta\bm{\mathcal{J}}_{\mathcal{X}}=-\iota_{\mathcal{X}}\left(\bm{E}\cdot\delta\phi\right)+\bm{\omega}\left(\phi;\delta\phi,\mathscr{L}_{\mathcal{X}}\phi\right)+d\left[\iota_{\mathcal{X}}\bm{\theta}\left(\phi;\delta\phi\right)\right],$ (40)
where it has not been assumed that $\phi$ satisfies the field equations nor that $\delta\phi$ satisfies the linearized field equations. Furthermore, it can be shown that $\bm{\mathcal{J}}_{\mathcal{X}}$ can be written in the form [23]
$\bm{\mathcal{J}}_{\mathcal{X}}=\bm{C}_{\mathcal{X}}+d\bm{Q}_{\mathcal{X}},$ (41)
where $\bm{Q}_{\mathcal{X}}$ is the Noether charge and $\bm{C}_{\mathcal{X}}\equiv\mathcal{X}^{a}\bm{C}_{a}$ with $\bm{C}_{a}=0$ being the constraint equations of the theory [31]. Having written in this form, we obtain the fundamental identity
$\bm{\omega}\left(\phi;\delta\phi,\mathscr{L}_{\mathcal{X}}\phi\right)=\iota_{\mathcal{X}}\bm{E}\cdot\delta\phi+\delta\bm{C}_{\mathcal{X}}+d\left[\delta\bm{Q}_{\mathcal{X}}-\iota_{\mathcal{X}}\bm{\theta}\left(\phi;\delta\phi\right)\right],$ (42)
It should be emphasized that this fundamental identity holds for arbitrary $\mathcal{X}^{a}$, $\phi$, and $\delta\phi$.
One immediate consequence of Eq. (42) is the gauge invariance of the symplectic form. If $\phi$ satisfies the equations of motion, $\bm{E}=0$, $\delta\phi$ satisfies the linearized constraints, $\delta\bm{C}_{a}=0$, and $\mathcal{X}^{a}$ is of compact support (or vanishes sufficiently rapidly at infinity and/or any boundaries), integration of Eq. (42) over a Cauchy surface $\Sigma$ yields
$W_{AB}\delta\phi^{A}\mathscr{L}_{\mathcal{X}}\phi^{B}=0.$ (43)
Consequently, the value of $W_{AB}\delta\phi^{A}\delta\phi^{B}$ is unchanged if either $\delta_{1}\phi$ or $\delta_{2}\phi$ is altered by a gauge transformation $\delta\phi\rightarrow\delta\phi+\mathscr{L}_{\mathcal{X}}\phi$ with $\mathcal{X}^{a}$ of compact support.
Another very important application concerns the case where $\mathcal{X}^{a}$ approaches a nontrivial asymptotic symmetry rather than being of compact support, in which case we can derive a formula for the Hamiltonian, $H_{\mathcal{X}}$, conjugate to the notion of “translations” defined by $\mathcal{X}^{a}$, and, thereby, a definition of ADM-type conserved quantities. Consider asymptotically flat spacetime with one asymptotically flat “end”. Integrating Eq. (42) over a Cauchy surface $\Sigma$, we have
$W_{AB}\delta\phi^{A}\mathcal{L}_{\mathcal{X}}\phi^{B}$
$=$ $\int_{\Sigma}\left(\iota_{\mathcal{X}}\bm{E}\cdot\delta\phi+\delta\bm{C}_{\mathcal{X}}\right)+\int_{S_{\infty}}\left[\delta\bm{Q}_{\mathcal{X}}-\iota_{\mathcal{X}}\bm{\theta}\left(\phi;\delta\phi\right)\right],$ (44)
@@
$\lim_{S\rightarrow S_{\infty}}\int_{S}\iota_{\mathcal{X}}\bm{\theta}\left(\phi;\delta\phi\right)=\lim_{S\rightarrow S_{\infty}}\delta\int_{S}\iota_{\mathcal{X}}\bm{B},$ (45)
@@
$W_{AB}\delta\phi^{A}\mathscr{L}_{\mathcal{X}}\phi^{B}=\delta H_{\mathcal{X}},$ (46)
@@
$H_{\mathcal{X}}=\int_{\Sigma}\bm{C}_{\mathcal{X}}+\int_{S_{\infty}}\left(\bm{Q}_{\mathcal{X}}-\iota_{\mathcal{X}}\bm{B}\right).$ (47)
@@
$\delta H_{\mathcal{X}}=\delta\phi^{A}D_{A}H_{\mathcal{X}}=\left(\mathcal{D}H_{\mathcal{X}}\right)_{A}\delta\phi^{A},$ (48)
@@
$W_{AB}\mathscr{L}_{\mathcal{X}}\phi^{B}=\left(\mathcal{D}H_{\mathcal{X}}\right)_{A}.$ (49)
@@
$\left(\mathscr{L}_{\mathcal{X}}\phi\right)^{A}=W^{AB}\left(\mathcal{D}H_{\mathcal{X}}\right)_{B},$ (50)
which is the usual form of Hamilton’s equations of motion on a symplectic manifold. Thus if both the asymptotic conditions on $\phi$ and the asymptotic behavior of $\mathcal{X}^{a}$ are such that a 3-form $\bm{B}$ satisfying Eq. (45) exists, then Eq. (47) yields a Hamiltonian conjugate to the notion of “translations” defined by $\mathcal{X}^{a}$. Note that when evaluated on solutions, $\bm{C}_{\mathcal{X}}=0$, $H_{\mathcal{X}}$ is purely a “surface term”
$H_{\mathcal{X}}|_{\bm{E}=0}=\int_{S_{\infty}}\left(\bm{Q}_{\mathcal{X}}-\iota_{\mathcal{X}}\bm{B}\right).$ (51)
In the case where $\mathcal{X}^{a}$ is asymptotic to a time translation $t^{a}$ at infinity, and $\bm{B}$ satisfying Eq. (45) can be found, then Eq. (51) defines the ADM mass
@@
In the case where $\mathcal{X}^{a}$ is asymptotic to a rotation $\varphi^{a}$ tangent to $\Sigma$ at infinity and $S$ is chosen so that $\mathcal{X}^{a}$ is tangent to $S$, the pull back of $\iota_{\mathcal{X}}\bm{\theta}$ to $S$ vanishes, then Eq. (51) with $\mathcal{X}^{a}=\varphi^{a}$ and $\bm{B}=0$ defines minus the ADM angular momentum